### v1

# On the emergence of numerical instabilities in Next Generation Reservoir Computing

Abstract.
Next Generation Reservoir Computing (NGRC) is a low-cost machine learning method for forecasting chaotic time series from data. However, ensuring the dynamical stability of NGRC models during autonomous prediction remains a challenge. In this work, we uncover a key connection between the numerical conditioning of the NGRC feature matrix — formed by polynomial evaluations on time-delay coordinates — and the long-term NGRC dynamics. Merging tools from numerical linear algebra and ergodic theory of dynamical systems, we systematically study how the feature matrix conditioning varies across hyperparameters. We demonstrate that the NGRC feature matrix tends to be ill-conditioned for short time lags and high-degree polynomials. Ill-conditioning amplifies sensitivity to training data perturbations, which can produce unstable NGRC dynamics. We evaluate the impact of different numerical algorithms (Cholesky, SVD, and LU) for solving the regularized least-squares problem.
Key words and phrases:
Reservoir Computing, Next Generation Reservoir Computing, Forecasting, Structured Matrices, Numerical Linear Algebra, Dynamical Systems
1. Lead Paragraph
Next Generation Reservoir Computing (NGRC) is a reservoir computing (RC) variant characterized by a reduced number of hyperparameters, data efficiency and forecasting quality, and easy interpretability. Despite its success in forecasting tasks, a central challenge remains: ensuring the dynamical stability of the model over long-term forecasts. Recent theoretical work points out a subtle interplay between approximation error and overfitting, yet no analytical method currently exists for selecting hyperparameters that guarantee the NGRC dynamical stability. Based on the premise that numerical instability during training can lead to dynamical instability, we investigate the connection between the choice of hyperparameters and the condition number of the NGRC feature matrix, formed by polynomial evaluations on time-delay coordinates. Notably, the NGRC model features structured matrices, such as Vandermonde-like and Hankel-like matrices, which have been extensively studied in numerical analysis. Merging numerical linear algebra and ergodic theory of dynamical systems, we characterize the conditioning of the feature matrix in terms of the maximum degree of the polynomial basis, the interplay between delay dimension and time lag, and the length of training data. This numerical analysis is a first step toward designing better strategies to select hyperparameters in reservoir computing.
2. Introduction
Forecasting time series is a fundamental problem across scientific disciplines [Cas89, MSW06, McE15], especially when dealing with chaotic systems time series [KS03]. In cases where the governing equations are unknown, the task becomes significantly more challenging. Reservoir Computing (RC) has emerged as a powerful machine learning method for forecasting chaotic dynamical systems directly from data [JH04, LPH+17, LHO18, PLH+17, PHG+18]. A recent promising variant, the Next Generation Reservoir Computing (NGRC) model, recasts the input data as a nonlinear vector autoregressive model [GBGB21]. It requires fewer hyperparameters, shorter training times, and less “warm-up” compared to traditional RC methods [BG22, ZC23]. NGRC stands out for its lower computational cost for forecasting but also other tasks, such as control [KBG24b], reconstruction from partial measurements [GBGB21, RP24], basin reconstruction [GFR22, ZC23] and experimental implementation [KBG24a, CMHR24, WHB+24].
Despite its success in forecasting chaotic time series, a central problem is ensuring the reservoir dynamical stability for short and long-term prediction [Luk12, LHO18, ZC23, GLO25]. Traditionally, in NGRC, training consists of solving a regularized least squares minimization problem (Tikhonov regularization or Ridge regression) to find the readout weight matrix that best predicts one step into the future. The trained model is then run autonomously for forecasting. Although NGRC might be successful during training, NGRC models might be unstable, diverging to infinity during the autonomous testing phase.
Recent theoretical work links reservoir dynamical stability to properties of the original dynamics and the approximation capacity used during training [STY98, DG00, BD23, BD23, GLO25]. By seeing reservoirs as high-dimensional embeddings [HHD20, GHO23], the largest Lyapunov exponent of the reservoir dynamics should not differ much from the original dynamics. However, there is no analytical method to guide hyperparameter selection that guarantees NGRC stability. Developing an analytical method for selecting hyperparameters requires understanding how these choices influence the NGRC dynamics. The mapping from hyperparameters to NGRC dynamics is often nontrivial. For instance, adjusting only the regularization strength [LHO18] can drastically alter the reservoir’s behavior. Moreover, an unstable NGRC model can emerge when the regularizer parameter is not scaled to the length of the training data [ZdSC25].
A promising direction comes from numerical linear algebra: for each hyperparameter configuration, we can examine the condition number of the feature matrix — the matrix that evaluates the input trajectory along the polynomial functions of time-delayed coordinates during training. The condition number quantifies the sensitivity of the least squares solution to perturbations [GVL96, TB97]. When the matrix is ill-conditioned (large condition number), even small changes in the data can produce large readout weights. These large weights are then fed back into the reservoir, potentially driving the model into dynamical instability [ZdSC25]. In this view, numerical instability during training (due to ill-conditioning) can translate into dynamical instability during prediction.
In this paper, we exploit a key structural insight of NGRC: its architecture naturally connects concepts from numerical linear algebra and dynamical systems theory. We show that the NGRC feature matrix possesses well-known structured forms — specifically, a Vandermonde-like matrix, due to the evaluation of polynomial basis functions on time series data, similarly done in polynomial interpolation; and a Hankel-like matrix, due to the use of time-delay coordinates. These matrix classes have been rigorously analyzed in numerical analysis, particularly in terms of their conditioning properties [GVL96, TB97, Hig02, BT17]. By leveraging this structure, we systematically characterize how the conditioning of the feature matrix varies with hyperparameter choices, and we pinpoint the regimes that lead to severe ill-conditioning — a potential precursor to NGRC instability. We use the Lorenz 63 system to demonstrate that the NGRC feature matrix becomes ill-conditioned, particularly for short time lags, high-degree polynomials, and moderate training lengths. Ill-conditioned matrices are a source of numerical instabilities, depending on the condition number and the residual error [GVL96, TB97, Hig02]. Here we test different numerical algorithms to solve the regularized least squares problem of the NGRC training: Cholesky decomposition, SVD, and LU decomposition. Our contribution is not to propose a new method, but to clarify why ill-conditioning arises in NGRC and how it affects long-term prediction stability.
This paper is organized as follows: the preliminaries are presented in Section 3.1. We then introduce the Lorenz system generated by the explicit forward Euler model as a toy model to motivate our problem formulation. Our main results on condition number characterization appear in Section 5. We follow this with numerical testing of different algorithms for solving the regularized least squares problem in Section 6. Finally, we summarize the results and present the discussion in Section 7.
3. Preliminaries
3.1. Notation
[…]
3.2. Learning dynamics
We assume a dynamical system exists $\mathbf{f}:M\to M$ lying on a compact $d-$dimensional metric space $M\subset\mathbb{R}^{d}$. For our purposes, $\mathbf{f}$ is given by discretizing an ordinary differential equation sampled uniformly at every $h$ time step. Instead of knowing $\mathbf{f}$, we only have access to a trajectory $\{\mathbf{x}_{n}\}_{n\geq 0}$ for a given initial condition $\mathbf{x}_{0}$, where $\mathbf{x}_{n}=\mathbf{f}^{n}(\mathbf{x}_{0})$, where $\mathbf{f}^{n}$ denotes the $n$-fold composition of the map $\mathbf{f}$. The problem is to learn the original dynamics $\mathbf{f}$ only from the observed data and be capable of forecasting it forward in time for unseen data.
3.3. Next Generation Reservoir Computing
We focus on using Next Generation Reservoir Computing (NGRC) [GBGB21], which recasts the input data as a nonlinear vector autoregression model [Bol21], akin to a numerical integration scheme. Differently from numerical integration schemes that aim to find the unknown $\{\mathbf{x}_{n}\}_{n\geq 0}$ from a known $\mathbf{f}$, NGRC aims to find an unknown $\mathbf{f}$ from a known $\{\mathbf{x}_{n}\}_{n\geq 0}$.
As the original formulation, the NGRC model utilizes a linear combination of polynomial functions evaluated at the current and time-delayed coordinates. More precisely, fix a delay dimension $k\in\mathbb{N}$ and time lag (or time skip) $\tau\in\mathbb{N}$, and define the embedding map $\mathbf{g}_{k,\tau}:M\to(\mathbb{R}^{d})^{k}$ by
(1) $\mathbf{g}_{k,\tau}(\mathbf{x})=\mathbf{vec}\Big{(}\mathbf{x},\mathbf{f}^{\tau}(\mathbf{x}),\dots,\mathbf{f}^{(k-1)\tau}(\mathbf{x})\Big{)},$
that is valid for any point $\mathbf{x}\in M$, and the coordinate in $(\mathbb{R}^{d})^{k}$ is denoted as $\mathbf{X}=\mathbf{g}_{k,\tau}(\mathbf{x})$. Specifically, for the time series $\{\mathbf{x}_{n}\}_{n\geq 0}$, the following holds:
$\mathbf{X}_{n}\equiv\mathbf{g}_{k,\tau}(\mathbf{x}_{n})$ $:=\mathbf{vec}\Big{(}\mathbf{x}_{n},\mathbf{f}^{\tau}(\mathbf{x}_{n}),\dots,\mathbf{f}^{(k-1)\tau}(\mathbf{x}_{n})\Big{)}$
$=\mathbf{vec}\Big{(}\mathbf{x}_{n},\mathbf{x}_{n+\tau},\dots,\mathbf{x}_{n+(k-1)\tau}\Big{)}.$
Let $\alpha\in(\mathbb{N}^{d})^{k}$ be the multi-index notation with $|\alpha|=\sum_{j=0}^{k-1}\sum_{i=1}^{d}\alpha_{i,j}$. The set of multivariate monomials in $kd$ variables up to degree $p$ is given by:
$\mathcal{P}_{p}^{kd}=\Big{\{}\mathbf{X}_{n}^{\alpha}=\prod_{j=0}^{k-1}\prod_{i=1}^{d}x_{n+j\tau,i}^{\alpha_{i,j}}:\quad|\alpha|\leq p\Big{\}},$
and has cardinality $m:=\binom{kd+p}{p}$. For example, if $k=1$, $d=2$, and $p=2$, then $\mathcal{P}_{2}^{2}$ includes terms like $x_{n,1}$, $x_{n,1}^{2}$, $x_{n,2}$, $x_{n,2}^{2}$, $x_{n,1}x_{n,2}$, and constant $1$. This induces a high-dimensional map $\psi:(\mathbb{R}^{d})^{k}\to\mathbb{R}^{m}$ as
(2) $\psi(\mathbf{X}_{n})=\mathbf{vec}\Big{(}\mathbf{X}_{n}^{\alpha}\Big{)}_{|\alpha|\leq p},$
which constructs a $m$-dimensional vector by the collection of monomials $\mathbf{X}_{n}^{\alpha}$.
Training phase. The NGRC model assumes that the time series can be recast akin as an iterative scheme induced by the high-dimensional map $\psi$
(3) $\mathbf{x}_{n+1}=\mathbf{x}_{n}+\mathbf{W}\psi(\mathbf{X}_{n-(k-1)\tau}),\quad n\geq(k-1)\tau.$
The goal of the training phase is to determine the readout matrix $\mathbf{W}\in\mathbb{R}^{d\times m}$ that best satisfies the above relation at each time step. The available time series $\{\mathbf{x}_{n}\}_{n\geq 0}^{N_{\mathrm{train}}+(k-1)\tau}$ has the number of points necessary to use Eq. 3, where $N_{\mathrm{train}}$ corresponds to the number of training data points and the warm-up $(k-1)\tau$ is the required number of data points to construct the embedded coordinate $\mathbf{X}_{n}$.
For the case of a sufficiently long time series ($N_{\mathrm{train}}>m$), finding $\mathbf{W}$ can be solved by a least squares with a Tikhonov regularization (or Ridge regression). Let us write explicitly the rows of $\mathbf{W}$, such that $\mathbf{W}=(\mathbf{w}_{1}^{\top},\dots,\mathbf{w}_{d}^{\top})$ where $\mathbf{w}_{i}$ is a $m$ dimensional vector. The regression can be made separately for each component of the input trajectory as follows: introduce for $i$-th component
(4) $\mathbf{y}_{i}=\begin{pmatrix}x_{(k-1)\tau+1,i}-x_{(k-1)\tau,i}\\ x_{(k-1)\tau+2,i}-x_{(k-1)\tau+1,i}\\ \vdots\\ x_{(k-1)\tau+N_{\mathrm{train}},i}-x_{(k-1)\tau+N_{\mathrm{train}}-1,i}\end{pmatrix}$
and construct the $N_{\mathrm{train}}\times m$ feature matrix using $\psi(\mathbf{X}_{n})$ as row vectors
(5) $\Psi=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\psi^{\top}(\mathbf{X}_{0})\\ \psi^{\top}(\mathbf{X}_{1})\\ \vdots\\ \psi^{\top}(\mathbf{X}_{N_{\mathrm{train}}-1})\end{pmatrix}.$
The solution to the following minimization problem yields each row of $\mathbf{W}$
(6) $\mathbf{w}_{i}(\beta)=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\Big{\{}\|\mathbf{y}_{i}-\Psi\mathbf{u}\|^{2}+\beta\|\mathbf{u}\|^{2}\Big{\}},$
where $\beta$ is the regularizer parameter. This is equivalent to solving $d$ independent regressions. The unique solution to this minimization problem is
(7) $\mathbf{w}_{i}(\beta)=(\Psi^{\top}\Psi+\beta\mathbf{1}_{m})^{-1}\Psi^{\top}\mathbf{y}_{i},$
where $\mathbf{1}_{m}$ is the $m\times m$ identity matrix.
Testing phase. The NGRC model then evolves autonomously according to the learned dynamics:
(8) $\mathbf{r}_{n+1}$ $=\mathbf{r}_{n}+\mathbf{W}\psi(\mathbf{R}_{n}),\quad n\geq 0,$
where $\mathbf{r}\in\mathbb{R}^{d}$ denotes the $d-$dimensional state variable of the NGRC model,
$\mathbf{R}_{n}=\mathbf{vec}\Big{(}\mathbf{r}_{n},\mathbf{r}_{n+\tau},\dots,\mathbf{r}_{n+(k-1)\tau}\Big{)},$
and the autonomous evolution is initialized using the last $(k-1)\tau$ values from the training set to construct the initial embedded coordinate given by the
$\mathbf{r}_{n}$ $=\mathbf{x}_{N_{\mathrm{train}}+n},\quad n=0,1,\dots,(k-1)\tau-1.$
The goal is that the NGRC trajectory $\{\mathbf{r}_{n}\}_{n\geq 0}$ predicts features from the unseen trajectory of the original dynamics. We denote the number of iterations during the testing phase as $N_{\mathrm{test}}$.
Figure 1. NGRC model accurately reproduces the Lorenz attractor. The top panel shows the NGRC reconstruction (red) of the Lorenz attractor (blue). The bottom panel shows the NGRC model’s performance over the training and testing phases. The horizontal axis shows time in Lyapunov time $\frac{1}{\Lambda}$ where $\Lambda=0.9056$ is the maximum Lyapunov exponent of the Lorenz system. The parameters are $h=0.01$, delay dimension $k=1$, maximum degree $p=2$, $N_{\mathrm{train}}=500$, $N_{\mathrm{test}}=10000$ (but a smaller time window is shown), and regularizer parameter $\beta=0$.
3.4. Key example: Lorenz system with Explicit forward Euler method
A widely studied chaotic system often used as a benchmark in forecasting tasks for machine learning algorithms is the Lorenz 63 system [Lor63]:
(9) $\begin{split}\dot{x}&=10(y-x)\\ \dot{y}&=x(28-z)-y\\ \dot{z}&=xy-\frac{8}{3}z.\end{split}$
Consider the time series given by integration of the Lorenz system vector field $\mathbf{F}$ using the explicit forward Euler method:
(10) $\mathbf{x}_{n+1}=\mathbf{x}_{n}+h\mathbf{F}(\mathbf{x}_{n}),$
which defines a discrete-time dynamical system that approximates the continuous flow of the original dynamics for small $h$. For this time series, let us consider the NGRC model with $k=1$ (absence of any time-delayed coordinates). Figure 1 shows the attractor reconstruction and forecasting capability of the NGRC model for the Lorenz system Eq. 9. The NGRC model captures the original dynamics, successfully predicting the trajectory for up to five Lyapunov times. This also extends to other dynamical features of the original dynamics.
E. Lorenz realized that extracting successive local maxima of the $z-$component yields a one-dimensional return map that is topologically conjugated to the tent map [Lor63], which is referred to as the successive local maxima map. The left panel in Fig. 2 shows that the NGRC model successfully captures this map. The right panel of Fig. 2 displays the power spectrum density of the three variables for the original dynamics and NGRC model, showing that the NGRC model also reproduces the long-term statistics of the Lorenz attractor. These observations demonstrate NGRC’s ability to recover the chaotic dynamics using only observed trajectory data, without explicit knowledge of the underlying equations.
Figure 2. NGRC model captures topological and statistical features of the original dynamics. The left panel displays the Poincaré return map of the successive local maxima of the original dynamics (in blue) and the reconstructed one (in red). The right panel depicts the power spectrum density of the original dynamics (in blue) and the reconstructed one (in red). Both curves mostly overlap for all frequencies, indicating that the NGRC model captures correctly the long-term statistics on the attractor. The parameters are the same used for Fig. 1.
3.5. Perturbation bounds for the least squares
This example is insightful when we examine the numerical accuracy of the least squares problem involved during the NGRC training. In this setup, the NGRC model trained with $k=1$ effectively learns an explicit forward Euler scheme, reducing to identifying the coefficients of a polynomial vector field from the time series data. For an arbitrary delay dimension $k$, using $\mathcal{P}_{p}^{kd}$ and the corresponding embedded vector $\mathbf{X}_{n}$, the explicit forward Euler discretization of a polynomial vector field implies that there exists a coefficient matrix $\mathbf{C}\in\mathbb{R}^{d\times m}$ such that:
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\mathbf{C}\psi(\mathbf{X}_{n-(k-1)\tau}),$
where $\mathbf{C}=(\mathbf{c}_{1}^{\top},\dots,\mathbf{c}_{d}^{\top})$ and each row vector $\mathbf{c}_{i}^{\top}$ corresponds to the coefficient vector of the $i$-th component of the vector field in the polynomial basis, where its magnitude depends on the time step $h$  [^1] Since the explicit forward Euler method evaluates the vector field at current state only, the true coefficients $\mathbf{c}_{i}$ have zero entries corresponding to monomials functions evaluated at any of the time-delayed coordinates.. Since the problem can be solved independently for each component $i$, let us temporally drop the $i-$th dependence. If $\Psi$ is of full column rank, then the least squares solution ($\beta=0$) is unique and is attained by $\mathbf{c}$:
$\mathbf{c}=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\big{\{}\|\mathbf{y}-\Psi\mathbf{u}\|^{2}\big{\}},$
where $\mathbf{y}$ is the vector in Eq. 4. However, in practice, numerical linear algebra is required to assess the accuracy of the computed solution. As a result, the NGRC training corresponds to solving a perturbed least squares problem of the form
(11) $\mathbf{w}=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\big{\{}\|\mathbf{y}+\Delta\mathbf{y}-(\Psi+\Delta\Psi)\mathbf{u}\|^{2}\big{\}},$
where $\Delta\mathbf{y}$ and $\Delta\Psi$ are small perturbation errors (e.g., due to rounding errors and discretization cumulative error). The sensitivity of the least squares problem describes how small perturbations in $\mathbf{y}$ and $\Psi$ affect the solution. The numerical accuracy in recovering $\mathbf{c}$ heavily depends on two factors:
  • The condition number of $\Psi$, $\kappa(\Psi)=\|\Psi\|_{2}\|\Psi^{\dagger}\|_{2}$, where $\Psi^{\dagger}$ is the pseudoinverse, or equivalently, $\kappa(\Psi):=\frac{\sigma_{1}}{\sigma_{m}}$ written in terms of the maximum and minimum singular values, $\sigma_{1}$ and $\sigma_{m}$, respectively.
  • The magnitude of the minimum residual $\|\mathbf{y}-\Psi\mathbf{w}\|$, measuring how well the model fits the data. This introduces the closeness of fit $\theta=\sin^{-1}\big{(}{\frac{\|\mathbf{y}-\Psi\mathbf{w}\|}{\|y\|}}\big{)}$. This angle characterizes the geometric misalignment between the $\mathbf{y}$ and the range of the feature matrix [TB97].
The following first-order bound quantifies numerical accuracy quantified by the relative error [GVL96, Theorem 5.3.1]: for $\varepsilon:=\max\Big{\{}\frac{\|\Delta\Psi\|}{\|\Psi\|},\frac{\|\Delta\mathbf{y}\|}{\|\mathbf{y}\|}\Big{\}}$ such that $\varepsilon\kappa(\Psi)<1$ and $\sin(\theta)\neq 1$:
(12) $\frac{\|\mathbf{w}-\mathbf{c}\|}{\|\mathbf{c}\|}\leq\varepsilon\Big{(}\frac{2\kappa(\Psi)}{\cos(\theta)}+\tan{\theta}\kappa(\Psi)^{2}\Big{)}+\mathcal{O}(\varepsilon^{2}).$ ) .
Even when $\Psi$ has full column rank, a large condition number (i.e., $\Psi$ is ill-conditioned) can cause the least squares solution to deviate drastically due to the perturbation. More precisely, $\varepsilon$ above can be quantified in terms of the machine precision, which in our case is $\varepsilon_{\mathrm{machine}}=\mathcal{O}(10^{-16})$. The heuristics is that if $\varepsilon_{\mathrm{machine}}=\mathcal{O}(10^{q})$ and $\kappa(\Psi)=\mathcal{O}(10^{q^{\prime}})$, the computed solution has about $q-q^{\prime}$ correct decimal digits. Hence, we say $\Psi$ is ill-conditioned with respect to the machine precision whenever $\varepsilon_{\mathrm{machine}}\kappa(\Psi)$ is large. For instance, if the condition number grows beyond $10^{13}$, the solution can have at best digits around $0.001$. The other quantity is the closeness of fit $\theta$. If the closeness of fit is large, the relative error might scale with the second term in Eq. 12, i.e., $\kappa(\Psi)^{2}$. Consequently, for an ill-conditioned $\Psi$, the numerical accuracy can be bad.
A similar result of Eq. 12 also holds for the Tikhonov regularization ($\beta>0$), where $\frac{\sigma_{1}}{\sqrt{\beta}}$ plays the role of the condition number [Han98, 5.1.1]. By increasing $\beta$, the condition number becomes smaller, and thus the regularized solutions are less sensitive to perturbations. However, increasing $\beta$ also makes the regularization error larger.
A well-established issue in numerical analysis [GVL96, TB97, Hig02] is the numerical stability for solving the least squares problem. To compare the impact of different numerical algorithms, we select standard methods in python using scipy [VGO+20] to solve the least squares problem, including:
  • Cholesky decomposition: The first method is based on solving the least squares via regularized normal equation for the unknown $\mathbf{u}$:
    (13) $(\Psi^{\top}\Psi+\beta\mathbf{1}_{m})\mathbf{u}=\Psi^{\top}\mathbf{y}_{i},\quad i=1,\dots,d,$
    which is computed numerically by Cholesky factorization in scipy.linalg.solve, assuming that the matrix is positive definite. The relative error for solving linear equations is slightly different from the least squares: it scales quadratically with the condition number $\kappa(\Psi)$, because of the factor $\Psi^{\top}\Psi$, and does not depend on $\theta$ (see [GVL96, TB97, Hig02] for detailed exposition). For $\beta>0$, we expect to scale as $\sigma_{1}^{2}/\beta$.
  • Singular Value Decomposition (SVD): The second method is written in terms SVD of $\Psi=\mathbf{U}\Sigma\mathbf{V}^{\top}$, which is computed via scipy.linalg.svd, with $\Sigma=\mathrm{diag}(\sigma_{1},\dots,\sigma_{m})$,
    (14) $\mathbf{w}_{i}=\Psi_{\beta}^{\dagger}\mathbf{y}_{i},$
    where $\Psi_{\beta}^{\dagger}:=\mathbf{V}\Sigma_{\beta}\mathbf{U}^{\top}$ is the regularized pseudo-inverse with the regularized singular values:
    $\Sigma_{\beta}=\mathrm{diag}\Big{(}\frac{\sigma_{1}}{\sigma_{1}^{2}+\beta},\dots,\frac{\sigma_{m}}{\sigma_{m}^{2}+\beta}\Big{)}.$
    Although this method can be computationally expensive, it is an option method to solve numerically the Tikhonov regularization [Han98]. Solving the least square problem with SVD is known to be stable [TB97, 19.4] [^2] The built-in implementation of SVD is based on a QR decomposition [ABB+99]. Solving least square problems based on QR decomposition satisfies Eq. 12. .
  • LU decomposition: The third method corresponds to solving Eq. 7 as it is mathematically written, where the inverse is computed numerically using LU decomposition via scipy.linalg.inv. This is similar to Choleskey decomposition, scaling quadractically with $\kappa(\Psi)$ for $\beta=0$ and as $\sigma_{1}^{2}/\beta$ for $\beta>0$.
Other commonly used numerical methods are pseudo-inverse and least square methods, scipy.linalg.pinv and scipy.linalg.lstsq, respectively, but those inherently add a built-in regularizer using the definition of $\varepsilon-$rank$(\Psi)$, where singular values are considered zero [GVL96, Theorem 2.5.3] whenever are upper bounded by $\sigma_{1}\times\varepsilon_{\mathrm{machine}}$.
The example of the Lorenz system generated by the explicit forward Euler method is also good in this regard. We expect that the closeness of fit is small because we would be fitting polynomial vector fields, using polynomial basis functions. As an illustration, for the parameters used in Fig. 1, the closeness of fit for these three algorithms is shown in Table 1. For this example, the condition number is $\kappa(\Psi)=8983.1469$. On the one hand, all results are of order $10^{-14}$, indicating that the fit is very close. On the other hand, there are differences among the algorithms. For this particular trajectory and parameters, SVD is the method with the minimum closeness of fit for all Lorenz’s components. As we will see in Section 6, SVD is the most robust method against the ill-conditioning of $\Psi$. For all simulations in this paper, we utilize the SVD algorithm, unless stated as in Section 6.
Algorithm $\theta_{x}$ $\theta_{y}$ $\theta_{z}$
Cholesky $3.12329\times 10^{-14}$ $8.88461\times 10^{-14}$ $1.32268\times 10^{-14}$
SVD $2.07528\times 10^{-14}$ $2.44480\times 10^{-14}$ $2.28179\times 10^{-15}$
LU $2.57541\times 10^{-14}$ $8.07677\times 10^{-14}$ $5.45502\times 10^{-14}$
Table 1. Resulting closeness of fit for Lorenz system, where $\theta_{x}$, $\theta_{y}$ and $\theta_{z}$ corresponds to the quantity for each component. The parameters are the same used for Fig. 1.
4. Problem statement
As aforementioned, the accuracy of a least squares solution is strongly affected by the conditioning of the feature matrix. Our focus here is to investigate under which conditions $\Psi$ becomes ill-conditioned. Specifically, we study how the condition number varies with the delay dimension $k$, the time lag $\tau$, and the length of the training data $N_{\mathrm{train}}$. We observe that $\Psi$ possesses rich internal structure: for $k>1$, it is simultaneously Vandermonde-like and Hankel-like because each entry corresponds to a monomial evaluated along a time-delayed trajectory of the original dynamics. As we will discuss, analyzing the conditioning of $\Psi$ is non-trivial and requires combining tools from numerical linear algebra with ergodic properties of dynamical systems.
Before proceeding, it is worth highlighting why this conditioning analysis matters. While the illustrative toy example in the previous section focuses on a polynomial vector field, it sheds light on learning dynamical systems from data using NGRC.
  • Governing equations from data and NGRC stability. When training NGRC, the task reduces to finding the governing equations of a dynamical system directly from time series data [CM87, BRT94a, BRT94b]. This is an intuitive perspective into the autonomous NGRC dynamics during testing. Even small modeling errors — especially for terms that are absent in the true system — can cause instability when fitting an ODE from data, as already identified in [YB07]. For example, consider we attempt to learn $\dot{x}=-x,\dot{y}=0$ but instead recover $\dot{x}=-x,\dot{y}=\varepsilon y$. Although $\varepsilon$ may be small, the trajectory of the learned system diverges due to exponential growth in $y$. Similarly, if $\Psi$ is ill-conditioned, the estimated readout weights may differ substantially from the true ones, leading to unstable NGRC dynamics [ZdSC25].
  • Regularization is necessary for ill-conditioned $\Psi$. Our analysis also clarifies when regularization ($\beta>0$) is required. Tikhonov regularization, as used in Eq. 6, is a standard method for mitigating the effects of ill-conditioned matrices. Selecting an appropriate $\beta$ is non-trivial and has been studied extensively [Han98]. For forecasting tasks, it has been a common approach to train the model to predict forward one step, and then select the regularizer parameter $\beta$ that ensures good prediction performance over multiple time steps ahead. Our results provide a theoretical foundation for this practice by identifying regimes where regularization is not just helpful but essential.
5. The structure of $\Psi$ characterizes its conditioning
We examine the matrix $\Psi$ in detail, observing that its structure resembles Vandermonde-like and Hankel-like matrices — objects that frequently arise in interpolation, signal processing, and numerical linear algebra [OK01, Ols01]. We numerically demonstrate that $\Psi$ can become ill-conditioned under various parameter settings. The underlying causes are analyzed separately: the influence of the maximum polynomial degree, the interplay between time lags and delay dimension, and the length of training data.
In all numerical experiments, the time series is generated using the explicit forward Euler method with step size $h=0.01$, discarding the initial transient time of $10000$ time steps. Moreover, the time series is normalized component-wise to be mapped inside the cube $[-1,1]^{3}$ by applying the transformation $\mathbf{x}_{n,i}\mapsto\frac{\mathbf{x}_{n,i}}{\max_{n}\{\mathbf{x}_{n,i}\}}$ for each component $i=1,\dots,d$. Since we are interested in the order of magnitude rather than the exact values of the condition number, we normalize $\Psi$ such that its columns have unit norm: for each column $\mathbf{u}_{i}$ of $\Psi$, the new columns are of the form $\frac{\mathbf{u}_{i}}{\|\mathbf{u}_{i}\|}$. This normalization is called column weighting [GVL96], which realizes a lower bound for the condition number of the original matrix [vdS69, Bec00].
5.1. Condition number grows exponentially with respect to $p$
Figure 3. Exponential growth wrt to maximum degree. Box plot of the condition number $\kappa(\hat{\Psi})$ over 25 distinct initial conditions for increasing maximum degree $p$. The exponential growth is confirmed by the exponential curve (black dashed line) plotted for reference. The numerical integrator is the explicit forward Euler method with step size $h=0.01$, and the number of training data points $N_{\mathrm{train}}=5000$.
The first numerical experiment increases the NGRC model’s maximum degree and evaluates the condition number $\kappa(\hat{\Psi})$ over different initial conditions. Fig. 3 shows that $\kappa(\hat{\Psi})$ increases exponentially with respect to $p$, with a growth rate approximately scaling as $e^{3p}$, as confirmed by the linear fit on the logarithmic scale. This suggests that, for systems requiring NGRC models with high-degree polynomials, the associated matrix $\Psi$ is unavoidably ill-conditioned.
The $\Psi$ structure can be related to a well-known matrix in numerical analysis and polynomial interpolation [Hig02]. Note that $\Psi$ evaluates multivariate polynomials along the time series, yielding a Vandermonde-like structure [KRS19]. For example, for $k=1$, $d=2$ and $p=2$:
$\Psi=\begin{pmatrix}1&x_{0,1}&x_{0,1}^{2}&x_{0,2}&x_{0,2}^{2}&x_{0,1}x_{0,2}\\ 1&x_{1,1}&x_{1,1}^{2}&x_{1,2}&x_{1,2}^{2}&x_{1,1}x_{1,2}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&x_{N_{\mathrm{train}}-1,1}&x_{N_{\mathrm{train}}-1,1}^{2}&x_{N_{\mathrm{train}}-1,2}&x_{N_{\mathrm{train}}-1,2}^{2}&x_{N_{\mathrm{train}}-1,1}x_{N_{\mathrm{train}}-1,2}\end{pmatrix}.$
There is a large body of evidence that Vandermonde matrices tend to be badly ill-conditioned [Pan16]. For the univariate case, the ill-conditioning is a consequence of the monomials being a poor basis for the polynomials on the real line. Monomials become increasingly highly correlated as the degree becomes larger over a finite interval. More specifically, it has been shown that the condition number of Vandermonde matrices for the univariate case has exponential growth with respect to $p$; see [GI87, KRS19]. In particular, the growth rate is at least $\big{(}\frac{2}{p+1}\big{)}^{1/2}(1+\sqrt{2})^{p-1}$ for any choice of real positive or symmetrically distributed points (nodes) [Bec00]; for other lower bounds estimates, see [Hig02]. Although precise results for the multivariate case (as relevant for NGRC) are lacking, our numerical observations are consistent with the exponential growth reported in the univariate case.
5.2. The interplay between time lags and delay dimension
Figure 4. Time lag makes $\Psi$ better conditioned. The left panel corresponds to the histogram of the condition number $\kappa(\hat{\Psi})$ over 100 distinct initial conditions for different combinations of NGRC models. Note that the $k=2$ (in orange), $k=3$ (in green), and $k=4$ (in red) overlap, showing a cluster of the different initial conditions at $10^{16}$. The right panel shows that the condition number monotonically decreases for larger time lag $\tau$. The solid line corresponds to the median, and the shaded areas correspond to $25\%$ and $75\%$ quantiles over 25 different initial conditions. The orange dashed line corresponds to the NGRC model with $k=1$, as shown in the left panel. The inset corresponds to the condition number of the submatrix $\hat{\Psi}_{x}$ for increasing $\tau$, whose decaying behavior is captured by the expression written for positive constant $K$ and exponent $\gamma$; see Example 5.1 for details. The numerical integrator is the explicit forward Euler method with step size $h=0.01$, maximum degree $p=2$, and the number of data points $N_{\mathrm{train}}=5000$.
Differently from the maximum degree $p$, which is a known issue in any polynomial interpolation using a monomial basis [Hig02], the effect of adding time-delayed coordinates and time lag is more specific to time-delayed models, such as the NGRC model. The numerical observation is that $\Psi$ is ill-conditioned for small $\tau$. The left panel of Fig. 4 shows the histogram of $\kappa(\hat{\Psi})$ over $100$ different initial conditions for different $k$. The case $k=1$ (in orange) corresponds to the NGRC model akin to the explicit forward Euler method — which will be the reference value (order $10^{2}$). As the delay dimension $k$ increases, the condition number abruptly shifts to $10^{16}$, indicating that the matrix has become rank deficient. We follow a similar argument of [ZdSC25] to pinpoint the specific reason:
Result. For small values of time lag $\tau$ and step size $h$, $\Psi$ is rank deficient for any delay dimension $k>1$.
Proof. Consider the case $\tau=1$. For $k>1$, there exists a subset of monomials in $\mathcal{P}_{p}^{kd}$ that evaluates the trajectory at consecutive time points, and their corresponding columns in $\Psi$ are linearly dependent on each other. This implies that the minimum singular value is close to zero, and consequently, $\Psi$ is rank deficient. For any univariate monomial in $\mathcal{P}_{p}^{kd}$, the argument is the following: choose two distinct columns of $\Psi$ corresponding to the same monomial evaluated at two consecutive time points. Let us denote such monomial as $\phi:\mathbb{R}^{d}\to\mathbb{R}$. By the explicit forward Euler method in Eq. 10 and $\phi$ being continuous differentiable function, the Mean-Value theorem implies that
[…]
where $\cdot$ corresponds to the usual inner product in $\mathbb{R}^{d}$. Two columns of $\Psi$ corresponding to $\phi$ evaluated along the time series for consecutive time points are given as
(16) $\mathbf{u}_{1}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\phi(\mathbf{x}_{0})\\ \phi(\mathbf{x}_{1})\\ \vdots\\ \phi(\mathbf{x}_{N_{\mathrm{train}}-1})\end{pmatrix}\quad\text{and}\quad\mathbf{u}_{2}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\phi(\mathbf{x}_{1})\\ \phi(\mathbf{x}_{2})\\ \vdots\\ \phi(\mathbf{x}_{N_{\mathrm{train}}})\end{pmatrix}.$
If any two columns of $\Psi$ are equal, it implies that $\Psi$ is singular, i.e., the minimum singular value $\sigma_{m}(\Psi)=0$. Consequently, by continuity of the singular values in terms of the entries of $\Psi$, $\mathbf{u}_{2}=\mathbf{u}_{1}+\mathcal{O}(h)$ with small $h$ implies that $\sigma_{m}(\Psi)$ is close to zero. A similar argument can be made for $\hat{\Psi}$, and other forward numerical integration schemes applied to the original differential equations. In agreement with the left panel of Fig. 4, the decreasing order of singular values of $\Psi$ has a clear gap, confirming that the matrix is rank deficient - results not shown.
We observe that the submatrices formed by evaluating the same monomial at different time-delays resemble a Hankel-like structure, which has already been analyzed in the context of data-driven dynamical systems in Dynamic mode decomposition [AM17]:
$\frac{1}{\sqrt{N_{\mathrm{train}}}}\begin{pmatrix}\phi(\mathbf{x}_{0})&\phi(\mathbf{x}_{1})&\dots&\phi(\mathbf{x}_{(k-1)\tau})\\ \phi(\mathbf{x}_{1})&\phi(\mathbf{x}_{1+\tau})&\dots&\phi(\mathbf{x}_{1+(k-1)\tau})\\ \vdots&\vdots&\ddots&\vdots\\ \phi(\mathbf{x}_{N_{\mathrm{train}}-1})&\phi(\mathbf{x}_{N_{\mathrm{train}}-1+\tau})&\dots&\phi(\mathbf{x}_{N_{\mathrm{train}}-1+(k-1)\tau})\end{pmatrix}.$
We utilize this structure to gain further insights on the influence of the time lag $\tau$ on the condition number, as we detail below.
5.2.1. Increasing the time lag improves the conditioning of $\Psi$
Although the exponential growth with respect to the maximum degree is inevitable, the ill-conditioning due to small $\tau$ vanishes for large time lags. For a fixed maximum degree and increasing the time lag makes $\Psi$ better conditioned, see the right panel of Fig. 4. It shows that for increasing the time lag, $\kappa(\hat{\Psi})$ decreases and asymptotically approaches the reference value (orange line), which corresponds to the condition number of the NGRC model with $k=1$ and $p=2$ as used in the left panel of Fig. 4. Increasing the time lag eliminates the spurious linear dependence between columns evaluating consecutive time points, which has also been observed in magnetic pendulum dynamics [ZdSC25].
The monotonic decay with respect to the time lag can be related to the statistical properties of the original dynamics: computing the condition number requires to compute the singular values of $\hat{\Psi}$, which are defined to be the square root of eigenvalues of $\hat{\Psi}^{\top}\hat{\Psi}$. First note that $\Psi^{\top}\Psi$ corresponds to the Euclidean inner product of columns $\mathbf{u}_{i}$ of $\Psi$, as in Eq. 16, where each entry is given by
(17) $\begin{split}(\Psi^{\top}\Psi)_{ij}=\langle\mathbf{u}_{i},\mathbf{u}_{j}\rangle&=\frac{1}{N_{\mathrm{train}}}\sum_{n=0}^{N_{\mathrm{train}}-1}\psi_{i}(\mathbf{X}_{n})\psi_{j}(\mathbf{X}_{n})\\ &=\frac{1}{N_{\mathrm{train}}}\sum_{n=0}^{N_{\mathrm{train}}-1}(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}(\mathbf{f}^{n}(\mathbf{x}_{0}))\\ &=:\frac{1}{N_{\mathrm{train}}}S_{N_{\mathrm{train}}}((\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau})(\mathbf{x}_{0}),\end{split}$
which corresponds to a Birkhoff average of the observable $(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}:M\to\mathbb{R}$. Then, the normalized version has the form $(\hat{\Psi}^{\top}\hat{\Psi})_{ij}=\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$, where $\mathbf{v}_{i}:=\frac{\mathbf{u}_{i}}{\|\mathbf{u}_{i}\|}$.
Since we are dealing with a chaotic system, it naturally calls for evoking its statistical properties. Let $\mu$ be an invariant ergodic measure preserved under $\mathbf{f}$. By the Birkhoff Ergodic theorem [VO16]: for any $\phi\in L^{1}(\mu)$:
[…]
For sufficiently chaotic dynamics, which is commonly captured by the decay of correlations over time, consider the case when the original dynamics is exponential mixing: $(\mathbf{f},\mu)$ is exponential mixing, such that the correlation function for any $\phi,\varphi,(\phi\cdot\varphi)\in L^{1}(\mu)$ decays:
[…]
For strong chaos, it is commonly associated with the exponential decay of correlations [Via97], including the Lorenz system 63 [AM16].
The Birkhoff Average theorem is valid for the observable $(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}$ if the pushforward probability measure $(\mathbf{g}_{k,\tau})_{*}\mu$ is an ergodic measure. This depends on the properties of the embedding map $\mathbf{g}_{k,\tau}$, which in turn depends on the delay dimension and time lag. Alternatively, since $\mathbf{g}_{k,\tau}$ evaluates monomial basis at different time, let us denote simply the monomial basis functions on the variables in $\mathbb{R}^{d}$ as $\phi_{i}:M\to\mathbb{R}$, i.e., $\phi_{i}\in\mathcal{P}_{p}^{d}$. Hence, the Birkhoff average in Eq. 17 has a particular form
(20) $\begin{split}\frac{1}{N_{\mathrm{train}}}S_{N_{\mathrm{train}}}((\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau})(\mathbf{x}_{0})=\frac{1}{N_{\mathrm{train}}}\sum_{n=0}^{N_{\mathrm{train}}-1}\phi_{i}\circ\mathbf{f}^{q}(\mathbf{f}^{n}(\mathbf{x}_{0}))\phi_{j}\circ\mathbf{f}^{q^{\prime}}(\mathbf{f}^{n}(\mathbf{x}_{0})),\qquad i,j=1,\dots,m,\\ q,q^{\prime}=0,1,\dots,\tau.\end{split}$
By invariance and ergodicity of $\mu$, it follows that for a large set of initial conditions, the Birkhoff average converges to
[…]
For an arbitrary $\mathbf{f}$ and the monomial basis $\mathcal{P}_{p}^{d}$, the right-hand side corresponds to the cross-correlation function for the observables $\phi_{i}$ and $\phi_{j}$ as in Eq. 19. The exponential decay of cross-correlations between delayed observables implies that the spurious linear dependencies among columns of $\Psi$ diminish as $\tau$ increases, as shown by the right panel in Fig. 4. For the monomial basis, it is a hard problem to be able to prove such a result. In contrast, we can study a toy case to grasp this intuition, see Example 5.1.
Example 5.1 ($x$-component case).
Consider the simplest case: $k=2$, $d=1$ and $p=1$, corresponding to $\mathcal{P}_{1}^{2}=\{1,x_{n},x_{n+\tau}\}$. We denote the submatrix formed only with these terms as $\hat{\Psi}_{x}$. Also, assume that the time series $\{x_{n}\}_{n\geq 0}$ has mean zero, which is approximately satisfied by the $x-$component of the Lorenz system due to the system’s symmetry. For this case, $\hat{\Psi}^{\top}\hat{\Psi}$ has the form
$\begin{pmatrix}1&0&0\\ 0&1&\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle\\ 0&\langle\mathbf{v}_{3},\mathbf{v}_{2}\rangle&1\\ \end{pmatrix},$
whose maximum and minimum eigenvalues are $1+\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle$ and $1-\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle$, respectively. Consequently, the condition number is given by
(21) $\begin{split}\kappa(\hat{\Psi})&=\sqrt{\frac{1+\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle}{1-\langle\mathbf{v}_{2},\mathbf{v}_{3}\rangle}}\\ &\leq\sqrt{\frac{1+Ke^{-\gamma\tau}}{1-Ke^{-\gamma\tau}}}=:h_{K,\gamma}(\tau),\end{split}$
for a positive constant $K$ and decay exponent $\gamma$, after applying the exponential decay of correlation bounds in Eq. 19. This formula $h_{K,\gamma}(\tau)$ shows that: while the function can be unbounded for small $\tau$, the function decays to one for large $\tau$. The inset of the right panel in Fig. 4 shows $\kappa(\hat{\Psi}_{x})$ (solid line) for increasing time lag, which decays monotonically to one. The dashed line corresponds to the fitted curve (with $K=1.01$ and $\gamma=0.01$) by the function $h_{K,\gamma}(\tau)$ in Eq. 21 to the numerical value. Interestingly, for $K=1$, the function can be written as $h_{1,\gamma}(\tau)=\coth^{\frac{1}{2}}{\big{(}\frac{\gamma\tau}{2}\big{)}}$.
We observe that the decay of $\kappa(\hat{\Psi})$ is slower than $\kappa(\hat{\Psi}_{x})$, which is related to the presence of the additional monomials. For instance, other observables as the $z-$component have slower correlation decay. A more detailed analysis of the full matrix will be left for future work.
Figure 5. Condition number with respect to the length of training data. Condition number of $\hat{\Psi}$ for increasing $N_{\mathrm{train}}$ for different time steps $h$. The left panel shows the results for the NGRC model with $k=1$ ($m=10$), and the right panel shows the model with $k=2$, $\tau=50$ ($m=28$). In both cases, the NGRC model has a maximum degree $p=2$. All curves for $h>0.1$ overlap for all $N_{\mathrm{train}}$ tested. The solid line corresponds to the median, and the shaded areas correspond to $25\%$ and $75\%$ quantiles over 10 different initial conditions.
5.3. Dependence on the length of training data
As we have seen, the statistical properties of the original dynamics give insights into the condition number of $\Psi$. This is not different from the length of training data $N_{\mathrm{train}}$. As aforementioned, to define the least squares problem during training, $\Psi$ should be tall and slim ($N_{\mathrm{train}}>m$); otherwise, the matrix has a non-trivial kernel and is rank deficient. To test the dependence on $N_{\mathrm{train}}$, Fig. 5 shows the condition number of $\Psi$ for $N_{\mathrm{train}}>\binom{kd+p}{p}$. For moderate values of $N_{\mathrm{train}}$, that is not much larger than $\binom{kd+p}{p}$, the condition number is not small. We observe that the condition number decays monotonically towards a limiting value $10^{2}$, but not uniformly. Initially, the decay is much faster than compared to larger values of $N_{\mathrm{train}}$. In the left panel, the NGRC model corresponds to the reference model we have been using so far, $k=1$ and $p=2$, and the right panel corresponds to the NGRC model with more terms, $k=2$, $\tau=50$, and $p=2$.
Since we are dealing with an ordinary differential equation, the length of the training data is determined by the time step $h$. Both panels of Fig. 5 show the dependence of $\kappa(\hat{\Psi})$ for different time steps. For this experiment, the numerical integration is performed using the explicit forward Euler method with time step $h=0.01$, then the time series is subsampled accordingly to the larger time steps. Similarly to $\tau$, the convergence to the limiting condition number is faster as the time step $h$ increases.
Notice that in both cases the condition number decays, but as expected, an NGRC model that contains more monomials requires more data points to reach the limit. In other words, the linear dependencies among columns can be long-lived as more monomials are used in the NGRC model. The convergence of $\kappa(\Psi)$ towards the limit and the difference between small and large $h$ can be explained due to the statistical properties of the Lorenz system, as we presented. The convergence follows from the Birkhoff averaging theorem in Eq. 18. For sufficiently long training data, the entries of $\hat{\Psi}^{\top}\Psi$ converge in a neighborhood of a limit value under different rates that depend on the pair of observables $(\psi_{i}\psi_{j})\circ\mathbf{g}_{k,\tau}$. This is in agreement with a recent result in traditional reservoir computing, which has shown that the Tikhonov regularization is an $L^{2}(\mu)$ approximation of ergodic dynamics [HHD21].
Similarly, $h$ in this particular construction plays a similar role as $\tau$ before. Larger $h$ corresponds to skipping the time evaluation of the monomial along time, and the exponential decay of correlations results in a smaller condition number. The fact that the entries of the matrix converge to a limit motivates scaling the matrix by the factor $\frac{1}{\sqrt{N_{\mathrm{train}}}}$. This is also relevant for the regularizer parameter, as we detail.
Remark 5.2 (Scaled regularizer parameter $\beta$ by the length of training data $N_{\mathrm{train}}$).
The scaling $\frac{1}{\sqrt{N_{\mathrm{train}}}}$ in the definition of $\Psi$ in Eq. 5 also scales $\beta$ with respect to the data length $N_{\mathrm{train}}$, which has been a common practice in statistics [GHW79], and more recently in reservoir computing [CAAG24, ZdSC25]. Let us denote $\tilde{\Psi}=\frac{1}{\sqrt{N_{\mathrm{train}}}}\Psi$. It follows that when $\tilde{\Psi}$ is used, the coefficient $\mathbf{w}(\beta)$ in Eq. 7 can be mapped to another coefficient with an adjusted regularizer parameter $N_{\mathrm{train}}\beta$. Note that
$\tilde{\mathbf{w}}(\beta)$ $=(\tilde{\Psi}^{\top}\tilde{\Psi}+\beta\mathbf{1}_{m})^{-1}\tilde{\Psi}^{\top}\mathbf{y}_{i}$
$=\frac{1}{\sqrt{N_{\mathrm{train}}}}(\frac{1}{N_{\mathrm{train}}}{\Psi}^{\top}\Psi+\beta\mathbf{1}_{m})^{-1}{\Psi}^{\top}\mathbf{y}_{i}$
$=\sqrt{N_{\mathrm{train}}}\mathbf{w}(N_{\mathrm{train}}\beta).$
6. Dynamical instability is algorithm-dependent
Figure 6. Ill-conditioning leads to numerical algorithm dependency. Comparison of three different numerical algorithms: Cholesky, SVD, and LU as the regularizer parameter $\beta$ is increased. Box plots in top to bottom panels as follows: valid prediction time (VPT), the distance between the induced successive local maxima maps, and the error $E$ between the power spectrum density, respectively. The colored hashed areas correspond to: absent/small (green), large (purple) regularization, and level of regularization in which all algorithms perform similarly (orange). The error bars are with respect to 50 different initial conditions. The orange band corresponds to the regularizer value in which all algorithms perform equally within statistical confidence. Whenever the algorithm failed to find a solution, we set $\mathbf{w}_{i}(\beta)=0$. Moreover, if the trajectory of the NGRC during testing was unbounded, we set the metric value to $-1$. This explains the blank spaces for Cholesky and LU decompositions. The parameters are $h=0.01$, delay dimension $k=2$, time lag $\tau=1$, the maximum degree $p=2$, and $N_{\mathrm{train}}=5000$.
As we have seen in Section 5, $\Psi$ can be ill-conditioned depending on the choices of maximum degree of polynomial, delay dimension, time lag, and length of training data. In this section, we consider the ill-conditioned regime for small $\tau$ and $h$, as commonly done in NGRC numerical experiments [GBGB21, GFR22, ZC23]. While Section 3.5 analyzed the algorithm’s accuracy under perturbations during NGRC training, here we focus on how such differences manifest in the NGRC’s resulting dynamics
This test mirrors a common practice in data-driven machine learning: fixing the hyperparameters, increasing the regularization parameter, and observing performance. This experiment highlights how different algorithms for solving the least-squares problem behave under varying regularization levels. We consider the NGRC model with $k=2$, $\tau=1$, $p=2$, and varying $\beta$ values, see horizontal axis of Fig. 6. The NGRC performance during testing is quantified by topological and statistical metrics, as described in detail in Section 8.1: valid prediction time (VPT) [VPH+20] quantifies for how long, in Lyapunov times, the NGRC trajectory is close to the original one; the distance between the induced successive local maxima map; and divergence error $E$ between the power spectrum densities estimated from the trajectories. Note that the columns of $\Psi$ are not normalized, as normalization can lead to different solutions [GVL96].
Fig. 6 shows the NGRC performance across different values of $\beta$ using the three algorithms. Starting from the right, the purple region corresponds to the regime of large $\beta$. The VPT decreases monotonically, where all three algorithms obtain a solution whose performance drops because the regularization is too strong. The distance between successive maps and the divergence error remain statistically indistinguishable across initial conditions, until they deteriorate due to excessive regularization.
The orange region corresponds to the regularizer value that all algorithms perform equally over all metrics. This regime corresponds to sufficient regularization for best performance in all three metrics. As opposed to smaller $\beta$, where the algorithms should overfit the input data, being susceptible to the influence of the ill-conditioning of $\Psi$. But this is not the case. The green regions show that for smaller $\beta$, the three algorithms yield different NGRC performances. Cholesky and LU decompositions exhibit discontinuous behavior as $\beta$ changes, whereas SVD varies smoothly. SVD is robust against the ill-conditioning of $\Psi$. Surprisingly, at $\beta=0$, its NGRC performance remains unaltered. Thus, regularization becomes unnecessary when an appropriate algorithm, such as SVD, is used. This result may have gone unnoticed, as the common practice is to apply sufficiently strong regularization to ensure good NGRC performance. And naturally, it points out to: we could attempt to minimize the condition number to avoid any conditioning-induced instabilities. In the next section, we show that minimizing the condition number does not guarantee dynamical stability.
6.1. Minimizing condition number does not imply dynamical stability
Figure 7. Better conditioning does not imply higher dynamical stability. The dots correspond to the median with respect to 25 different initial conditions, and error bars are $25\%$ (lower bound) and $75\%$ (upper bound) quantiles. The orange band represents the $\tau=15$, which shows the NGRC model with best performance in both metrics, VPT and divergence error $E$, but condition number $\kappa(\Psi)$ at $10^{4}$, which is not the minimum condition number possible with respect to $\tau$. If the trajectory of the NGRC during testing was unbounded, we set the metric value to $-1$. The parameters are $h=0.01$, delay dimension $k=2$, the maximum degree $p=5$, $N_{\mathrm{train}}=10000$ and regularizer parameter $\beta=0$.
In this section, we discuss an important observation: better conditioning of the feature matrix does not guarantee dynamical stability. For this test, we consider a partial measurement of the Lorenz system, having access only to the $x-$component time series $\{x_{n}\}_{n\geq 0}$, which is normalized to be in $[-1,1]$ as previous sections, see the left panel of Fig. 8. Recently, it has been shown that NGRC can forecast the $x-$component time series [RP24]. The underlying dynamics is not the Euler discretization of a polynomial vector field, but an unknown function of the current time and time-delayed coordinates by Takens’ embedding theorem [PCFS80, Tak81]. Thus, the unknown function does not necessarily lie in the span of $\mathcal{P}_{p}^{kd}$ and must be approximated by them, a scenario more typical in NGRC applications, and distinct from earlier sections.
The NGRC model has a maximum degree $p=5$ and $k=3$, which is chosen based on the standard false nearest neighbors method for estimating embedding delay dimension [KS03]. For the time lag, we proceed as follows: we increase the $\tau$ and measure the NGRC performance by the valid prediction time and divergence error between the power spectrum densities, see Fig. 7.
As expected by our analysis in Section 5.2, increasing $\tau$ reduces the condition number monotonically. However, NGRC performance exhibits a different trend. At first, the NGRC performance has a peak at $\tau=15$ highlighted by the orange band. But then, the performance worsens. This agrees with former observations in constructing embeddings of partial measurements of chaotic systems. For small $\tau$, the $x-$component time series is highly correlated with each other, while for large $\tau$, the time series become highly uncorrelated. In both cases, capturing the embedding becomes challenging [FS86, KS03]. The good time lag $\tau$ corresponds to the first minimum of mutual information. Interestingly, the NGRC performance has a peak and coincides with the time lag that attains the first minimum of mutual information. This illustrates that although better conditioning ensures numerical stability during training, it does not guarantee dynamical stability in NGRC predictions. An example of the NGRC trajectory and power spectrum at $\tau=15$ is depicted in Fig. 8.
Figure 8. NGRC learns partial measurements of the Lorenz system. The left panel shows the NGRC trajectory (red) compared to the original (blue). The right panel shows the comparison between the power spectrum densities. NGRC The parameters are $h=0.01$, delay dimension $k=2$, time lag $\tau=15$, the maximum degree $p=5$, $N_{\mathrm{train}}=10000$ and regularizer parameter $\beta=0$.
7. Discussion and conclusions
This work has presented a detailed numerical analysis of the feature matrix conditioning in NGRC. By leveraging the Vandermonde- and Hankel-like structures of the feature matrix and the statistical properties of the Lorenz system, we elucidated the interplay between hyperparameters and feature matrix conditioning. Through numerical experiments, we demonstrated that the condition number of $\Psi$ can become prohibitively large: (i) high maximum degree $p$, (ii) short time delay $\tau$, and (iii) moderate length of training data $N_{\mathrm{train}}$. We also observed that the choice of numerical algorithm used for training substantially impacts robustness to ill-conditioning.
Incidentally, in the presence of an ill-conditioned feature matrix, SVD-based training can lead to a stable NGRC even without regularization. This dependence on the numerical method suggests a practical step: evaluate the conditioning of the feature matrix and prefer SVD-based training before deciding whether regularization is necessary. While SVD may be computationally expensive for large systems, alternatives for efficient Tikhonov regularization exist, including classical approaches such as Élden’s method [Eld77] and modern techniques for large-scale inverse problems [LRV12].
Interestingly, while bad conditioning may cause NGRC instability, it does not imply it. For instance, we showed that a successful reconstruction of the $x$-component dynamics remains possible despite a non-minimal condition number of $\Psi$. This correspondence of conditioning and NGRC stability suggests that minimizing the condition number is not a universally optimal objective. In fact, classical numerical analysis states that conditioning is highly sensitive to the distribution of points being evaluated (collocation points or nodes) and the choice of polynomial functions [KRS19]. For example, Vandermonde matrices evaluated at points nearly equally spaced on the unit circle or using orthogonal polynomial bases (like Chebyshev) are known to be well-conditioned. These properties, however, are unrelated to the underlying dynamics. This observation raises an exciting research direction of constructing a basis of orthogonal polynomials adapted to the distribution of points generated by the trajectory of the dynamical system. A first step towards this direction has been applied to coupled maps for network reconstruction [PdSvS23].
One of the primary contributors to ill-conditioning is the use of high-degree polynomials. A natural strategy to mitigate this is to select a subset of monomials most relevant to the dynamics. However, identifying the optimal combination of terms is computationally challenging [DMA97]. Greedy algorithms, though suboptimal, have been successfully applied in reservoir computing [DSV+09]. More sophisticated approaches might impose prior structural information, such as sparsity or symmetry, aligning with recent developments in physics-informed machine learning and reservoir computing [CLS21, GGNFN23, GGRG24, BGR+21]. Developing practical algorithms to construct NGRC models that preserve or approximate selected features of the original system remains an open direction.
For complex dynamics, long polynomial expansions (large $p$) may be unavoidable, resulting in severely ill-conditioned matrices and expensive computations. In such cases, one must balance accuracy, interpretability, and computational cost. Recent alternatives address these challenges: for interpretable models, local low-degree polynomial approximations have been proposed [GPB25]; for efficiency, recursive Volterra series expansions offer a delay-agnostic, lower-cost alternative to NGRC [GTO25].
Finally, while this study focused on NGRC, the insights presented here are broadly relevant to other data-driven modeling approaches relying on solving linear equations and least-squares problems. In particular, our findings resonate with similar concerns in Dynamic Mode Decomposition (DMD) [DMM19, PD20], where matrix conditioning also plays a central role. Our results highlight the need for careful hyperparameter selection to ensure numerical stability of NGRC models. This numerical analysis is a first step toward designing better strategies to construct robust reservoir computers.
8. Appendix
8.1. Metrics
The reconstructed dynamics of the NGRC model is compared to the original dynamics using different metrics as detailed below.
Valid prediction time. We use the valid prediction time [VPH+20] to quantify the number of iterations such that the trajectories of the NGRC model and the original dynamics remain close to each other under a certain threshold. For a fixed threshold $\eta=0.9$, we define the valid prediction time as
(22) $\begin{split}n_{VPT}&:=\min_{n}\Big{\{}n:\frac{\|\mathbf{x}_{n}-\mathbf{r}_{n}\|}{\sqrt{\frac{1}{N}\sum_{k=0}^{N}\|\mathbf{x}_{k}\|^{2}}}>\eta\Big{\}}\\ VPT&:=\frac{1}{T_{\Lambda}}n_{VPT}h,\end{split}$
where $T_{\Lambda}=\frac{1}{\Lambda}$ is the maximum Lyapunov exponent of the original dynamical system. For instance, for the Lorenz system, $\Lambda=0.9056$.
Distance between successive maxima map. We introduce a distance between the induced successive maxima map of the original dynamics and the NGRC model. To this end, we utilize the successive local maxima from the time series of $z$-coordinate of the original and the NGRC model, $\{z_{n}^{\max}\}_{n\geq 0}$ and $\{\hat{z}_{n}^{\max}\}_{n\geq 0}$, respectively. To compute a smooth approximation of the successive maxima maps:
$S:z_{n}^{\max}\mapsto z_{n+1}^{\max},\quad\hat{S}:\hat{z}_{n}^{\max}\mapsto\hat{z}_{n+1}^{\max},$
we use B-spline interpolation of degree $3$. Let $\mathcal{I}\subset\mathbb{R}$ be the overlapping domain of both maps, then we define the mean absolute difference over $1000$ evenly spaced points $\{u_{j}\}_{j=1}^{N}\subset\mathcal{I}$ as:
$D=\frac{1}{1000}\sum_{j=1}^{1000}|S(u_{j})-\hat{S}(u_{j})|.$
Power spectrum density. To capture the long time statistics of the trajectories we computed the power spectrum density. In the case the trajectories sample the attractor similarly over time, the power spectrum density should be close to other. The attractor similarity is computed using Kullback-Leibler (KL) divergence between the two empirical power spectral densities of the trajectories. More precisely, from the trajectories $\{\mathbf{x}_{n}\}_{n}$ and $\{\mathbf{r}_{n}\}_{n}$ we estimate the power spectrum density $P_{i}$ and $\hat{P}_{i}$ of each $i$th component, respectively, using Welch’s method [Wel67] with Hann window, and compute the divergence error $E$ based on the KL divergence
$D_{KL}(P_{i}||\hat{P}_{i})$ $=\sum_{p=0}^{\frac{L}{2}+1}P_{i}(f_{p})\log_{10}\Big{(}\frac{P_{i}(f_{p})}{\hat{P}_{i}(f_{p})}\Big{)},$
$E$ $=\sum_{i=1}^{d}D_{KL}(P_{i}||\hat{P}_{i}).$
where $L=\lfloor\frac{5}{dt}\rfloor$ the number of segments, the number of overlapping points $L/2$, computing the periodograms with mean.
Author Declarations
Acknowledgments. The authors thank Zheng Bian for enlightening discussions. ERS and EB thank the Collaborative Research in Computational Neuroscience (CRCNS) through R01-AA029926. The ONR, ARO, NIH, and DARPA also support EB.
Conflict of Interest. The authors have no conflicts to disclose.
Author Contributions. ERS: Conceptualization (equal); Formal analysis (lead); Investigation (lead); Methodology (lead); Software (lead); Visualization (lead); Writing – original draft (lead); Writing– review & editing (equal). EB: Conceptualization (equal); Supervision (lead); Writing– review & editing (equal); Funding acquisition (lead); Project administration (lead);
Data Availability Statement
The code to produce the data that support the findings of this study is available in the https://github.com/edmilson-roque-santos/numerical_stability_NGRC.